Circular No. 8479

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005ab AND 2005ad Further to IAUC 8478, S. Nakano reports the discovry by K. Itagaki of

an apparent supernova (red mag 17.4) on numerous unfiltered CCD frames taken around Feb. 6.402 UT; Itagaki reports the new object at mag 16.6 on a CCD image taken on Feb. 7.434. SN 2005ad is located at α = 2^{h}28^{m}29^{s}.45, δ = −1^{o}08′20′′.0 (equinox 2000.0), which is 23′′ east and 45′′ north of the center of NGC 941. Nothing is visible at this location on CCD frames taken by Itagaki on Jan. 8 (limiting mag 19) and on 2003 Sept. 28 and earlier (limiting mag 19.5); it is also absent from an unspecified Digitized Sky Survey image.

Further to IAUC 8478, Nakano reports that Itagaki found SN 2005ab at mag 16.8 on a CCD image taken on Feb. 6.644 UT.

SUPERNOVA 2005W IN NGC 691 N. Elias-Rosa, H. Navasardyan, A. Harutunyan, S. Benetti, and M.

Turatto, Osservatorio Astronomico di Padova; A. Pastorello, Max-Planck-Institut; and F. Patat, European Southern Observatory, on behalf of the European RTN (IAUC 7987), report that a spectrogram of SN 2005W (cf. IAUC 8475), obtained on Feb. 2.7 UT with the Asiago 1.8-m telescope (+ AFOSC; range 355–780 nm, resolution 2.4 nm), shows it to be a type-Ia supernova about a week before maximum. The spectrum closely resembles that of SN 2002bo (Benetti et al. 2004, MNRAS 348, 261) at comparable phase. The expansion velocity deduced from the Si II 635.5-nm minimum is ∼ 11600 km/s (adopting the NED recession velocity of 2665 km/s for NGC 691).

COMET C/1995 O1 (HALE-BOPP) A. Rivkin and R. Binzel, Massachusetts Institute of Technology, report

that images of C/1995 O1 obtained on Jan. 8 with the Magellan Observatory’s Clay 6.5-m telescope (+ SDSS g′, r′, and i′ filters) show a tail at least 8′′.5 long (through g′ and r′ filters) and the following coma magnitudes in a 4′′.2 aperture: g′ = 20.73, r′ = 20.33, i′ = 20.06.

NOVA IN M31 Additional R-band magnitudes by K. Hornoch of the nova announced

on IAUC 8461: Jan. 11.719 UT, 15.16; 13.905, 15.04; 15.702, 15.27; 16.900, 15.34; 19.827, 15.50; 23.734, 15.80; 27.774, 16.61; 30.712, 17.17.

2005 February 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green